OpenStudy (anonymous):

Find the probability. An IRS auditor randomly selects 3 tax returns (without replacement) from 47 returns of which 5 contain errors. What is the probability that she selects none of those containing errors? A. 0.0012 B. 0.0006 C. 0.708 D. 0.7136

OpenStudy (amistre64):

i was thinking the hypergeometric probability

OpenStudy (amistre64):

\[P(x)=\frac{\binom{N-M}{n-k}\binom{M}{k}}{\binom{N}{n}}\] i think is how its set up

OpenStudy (amistre64):

N items M successes n drawn k of them are success

OpenStudy (amistre64):

n= 3 N=47 M=47-5 What is the probability that she selects none of those containing errors (3) successes is how im interping it

OpenStudy (amistre64):

(5 choose 0)(42*41*40/3!)\(47*46*45/3!); thnx to wolfram = .7079 if its right :)